OpenStudy (anonymous):

Which related equation can be used to solve the equation? Column A Column B 1. m + 5 = –7.3 2. m + 5 = 7.3 3. m + 7.3 = –5 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (texaschic101):

1. m + 5 = -7.3 m = -7.3 - 5 2. m + 5 = 7.3 m = 7.3 - 5 3. m + 7.3 = -5 m = -7.3 - 5 4. m + 7.3 = 5 m = -7.3 + 5